OpenStudy (anonymous):

Simplify using common denominator: a) (sin^3theta/cos theta) - (cos^3/sin theta) b) (cos^2theta/tan theta) - (sin^2theta/cot theta) c) (sec theta/cos theta) + (cosec theta/sin theta) d) [(costheta . sintheta)/cot theta] + [(costheta . sintheta)/tan theta]

OpenStudy (anonymous):

Srry for a really long question

OpenStudy (anonymous):

Add a medal if you want me to draw it

OpenStudy (confusionist):

...You mean you want us to give you a medal so you add more to your question? You shouldnt be asking questions to receive medals...

OpenStudy (anonymous):

\[ \frac{\sin^3(x)}{\cos(x)} - \frac{\cos^3(x)}{\sin(x)}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

Im srry @Confusionist i am new to this 'openstudy' thing. :)

OpenStudy (anonymous):

@satellite73 how do you actually do the question

OpenStudy (confusionist):

Nah, it's cool! You actually give a medal to the person that you think gave the best response by clicking 'best response' next to their name. :) Why don't you give it a try? I can help you out if you need anything else on the site. If you really like someone's answer you can become their fan and even give them a testimonial.

OpenStudy (anonymous):

Alright thanks @Confusionist

OpenStudy (mertsj):

This is really 4 questions, right?

OpenStudy (anonymous):

Yes... im sorry for bothering all of you!

OpenStudy (mertsj):

Is the first one what Satellite posted?

OpenStudy (anonymous):

take them 1 at a time

OpenStudy (mertsj):

Is the first one what Satellite posted?

OpenStudy (anonymous):

yes but i dont know how to finish it..

OpenStudy (mertsj):

The instructions say to get a common denominator which would be sin(x)cos(x)

OpenStudy (mertsj):

\[\frac{\sin^3x}{\cos x}\times \frac{\sin x}{\sin x}-\frac{\cos ^3x}{\sin x}\times\frac{\cos x}{\cos x}\]

OpenStudy (mertsj):

\[\frac{\sin ^4x-\cos ^4x}{\sin x \cos }=\frac{(\sin ^2x-\cos ^2x)(\sin ^2x+\cos ^2x)}{\sin x \cos x}=\frac{\sin ^2x-\cos ^2x}{\sin x \cos x}\]

OpenStudy (anonymous):

Wow Thanks! Very nice of you to write this all down! :)

OpenStudy (mertsj):

Now you try the next one.

OpenStudy (anonymous):

ok thats right. I should learn from this

OpenStudy (anonymous):

btw srry if i dont know/clueless to all this work because this is extension for me (im in yr 10 in a school in Australia)

OpenStudy (mertsj):

\[\frac{\cos ^2x}{\tan x}-\frac{\sin ^2x}{\cot x}=\frac{\cos ^2x}{\tan x}\times\frac{\cot x}{\cot x}-\frac{\sin ^2x}{\cot x}\times\frac{\tan x}{\tan x}=\]

OpenStudy (mertsj):

\[\frac{\cos ^2x \cot x}{\tan x \cot x}-\frac{\sin ^2x \tan x}{\tan x \cot x}=\frac{\cos ^2x \cot x-\sin ^2xtanx}{\tan x \cot x}\]

OpenStudy (mertsj):

\[\cos ^2x \times\frac{\cos x}{\sin x }-\sin ^2x \times\frac{\sin x}{\cos x}=\frac{\cos ^3x}{\sin x}-\frac{\sin ^3x}{\cos x}\]

OpenStudy (mertsj):

\[\frac{\cos ^4x-\sin ^4x}{\sin x \cos x}\]

OpenStudy (mertsj):

\[\frac{(\cos ^2x-\sin ^2x)(\cos ^2x+\sin ^2x)}{\sin x \cos x}= \frac{\cos ^2x-\sin ^2x}{\sin x \cos x}\]